For the following function, C computes the cost in millions of dollars of implementing a city recycling project when x percent of the citizens participate. c(x)=(1.3x)/(100-x)
a)Using this model, approximately what is the cost if 40% of the citizens participate? Include units with your answer
b)Using this model, estimate the percentage of participation that can be expected if $3 million is spent on this recycling project? Set up an equation and solve algebraically. Round to the nearest whole percent.

Answer by stanbon(75887)

For the following function, C computes the cost in millions of dollars of implementing a city recycling project when x percent of the citizens participate. c(x)=(1.3x)/(100-x)
a)Using this model, approximately what is the cost if 40% of the citizens participate? Include units with your answer
c(40) = (1.3*40)/(100-40)
c(40) = 52/60
c(40) = 0.866667 million dollars
-------------------------------------

b)Using this model, estimate the percentage of participation that can be expected if $3 million is spent on this recycling project? Set up an equation and solve algebraically. Round to the nearest whole percent.
3 million = (1,3x)/(100-x)
300 - 3x = 1.3x
300 = 4.3x
x = 69.77%
=========================
Cheers,
Stan H.